[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Jul 28, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Jul 28 12:13:26 UTC 2017


Here is today's summary of NHC computer operations:
--- PS

----------

TAFB:
1.) Issues since accessing ATCF from TAFB yesterday. M. Sardi/Doug Gear  
dealt with it. Related to NetApp migration.

2.) After successfully receiving the 12Z EPAC, 18Z ATL, and 00Z EPAC  
yesterday in AWIPS the 06Z ATL never came back. It appears overnight a  
second run was submitted before the prev one completed causing problems.

We are looking at this now (C. Mello and D. Gaer) to reset things.  
Meanwhile, I have asked C. Mello to look for a better way to give TAFB a  
way to monitor the status of a current run before attempting to lunch  
another. More to follow from Chris on this.


HSU:
1.) One CAVE session on the AWIPS-2 at the JHT desk refused to respond to  
the pan/zoom of the mouse. Another CAVE session on the same workstation had  
no such problem. I will have Chris Mello follow this up. We need to stress  
to the staff the need for them to open a ticket on the spot when it happens  
to given NCF the chance to collect logs and troubleshoot these problems.

2.) FSSE not coming at all. Monica is already working this problem.
Other items:

NHC RFCs:
1.) NetApp Migration continues today with big switch taking place today.  
Will need to monitor carefully over the weekend for any potential problems.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170728/2488d4ad/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list